Leg Spring (Running)

Running without springs would be exhausting. Consider a 70 kg person running at 4 m/sec. During each step their COM sinks 4 cm (below resting) during stance and rises 2 cm (above resting) during the aerial phase.

  1. Determine the positive work they must do against gravity with each step. Then, calculate the power they must put out with each step.

  1. If their forward speed fluctuates by 10% during each stance phase (i.e., 3.6 m/s at mid-stance), determine the work they must do to increase their velocity (i.e. the change in kinetic energy).
Answer #1

Work done = force *displacement

=70kg* 4m/sec ^2* (4-2)/100m ( average duration for one gait cycle is 0.98 to 1.08second, so to ease our calculation we will take 1 sec approx time to complete one gait cycle)

= 70*4*2/100= 70*4/50 = 7*4/5 = 28/5 = 5.6 Nm^2/sec^2.

Power= work/ time

= 5.6/1= 5.6 watt